OpenStudy (anonymous):

Simplify the expression: (1-cosx)(1+cosx)/sinx

OpenStudy (anonymous):

set a = 1 and b = cos(x) then the expression becomes: \[\frac{ (a-b)(a+b) }{ \sin(x) }\] which equals: \[\frac{ a^2-b^2 }{ \sin(x) }\] now plug back in a and b: \[\frac{ 1-\cos^2(x) }{ \sin(x) }\] now you know the identity: \[\cos^2(x)+\sin^2(x)=1\] so:\[1-\cos^2(x)=\sin^2(x)\] so if you plug that in to your expression you get: \[\frac{ \sin^2(x) }{ \sin(x) }\] which simply equals:\[\sin(x)\] ANSWER: \[\sin(x)\]